OpenStudy (lopezking1):

3over4x+4=10 show work

OpenStudy (jazzyfa30):

is it written like this\[\frac{ 3 }{ 4 }x+4=10\]

OpenStudy (lopezking1):

Yes @jazzyfa30

OpenStudy (jazzyfa30):

ok so what do you do first? do you know?

OpenStudy (anonymous):

u should take whole numbers to one side 3/4 X =10-4 3/4 X =6

OpenStudy (lopezking1):

would you simplify the fraction into a decimal

OpenStudy (anonymous):

if ur're using a calculator ...its fine...but otherwise leave it in frac form

OpenStudy (jazzyfa30):

no just multiply the denominator by both sides

OpenStudy (lopezking1):

Oh yeah lol xD

OpenStudy (lopezking1):

12x=24?

OpenStudy (jazzyfa30):

now divide

OpenStudy (lopezking1):

x=12

OpenStudy (jazzyfa30):

no not quite what do you multiply 12 by to get 24

OpenStudy (lopezking1):

oh no i mean 2

OpenStudy (jazzyfa30):

exactly you got it

OpenStudy (anonymous):

actually ur kinda wrong cos..... 6 divded by 3/4 =8<<thats the answer

OpenStudy (jazzyfa30):

no it ain't I know what I'm talking about look at the question again

OpenStudy (anonymous):

ok substitute x=8 in ur 1st comment to the question

OpenStudy (lopezking1):

No yeah @physicsboffin is right i did it on the calculator lol

OpenStudy (jazzyfa30):

oh ok

OpenStudy (anonymous):

yep im not tryin to be rude or anything @jazzyfa30 ...

OpenStudy (lopezking1):

so the answer is x=8

OpenStudy (anonymous):

yep

OpenStudy (lopezking1):

coolios

OpenStudy (anonymous):

encantados de ayudarle